# Analysis of Semi-Supervised Learning on Hypergraphs

### v1

@@ 4.3 Higher-order hypergraph learning
For our first result, the proof follows from an application of [90, Proposition 4.21]. The key observation is that for any $\varepsilon_{n}$, $p>0$ and $v\in\mathrm{C}^{\infty}(\bar{\Omega})$ with $v(x_{i})=y_{i}$ for $i\leq N$, we can pick $\{(\nu_{n},v_{n})\}_{n=1}^{\infty}=\{(\mu_{n},v|_{\Omega_{n}})\}_{n=1}^{\infty}$ and hence the same recovery sequence for $\mathcal{J}^{(p_{k})}_{n,\Delta_{n,\varepsilon_{n}^{(k)}}}$ with $1\leq k\leq q$. This allows us to use the subadditivity of $\limsup$ to deduce the result.
@@
The next result is shown analogously to Proposition 4.15. In particular, one relies on the compactness result [90, Proposition 4.13]: we only require that our smallest length-scale $\varepsilon_{n}^{(q)}$ satisfies the appropriate upper bound and that the its associated power $p_{q}$ scales correctly with the dimension of $\Omega$. Then, the problem reduces to using the superadditivity of $\liminf$ and [90, Theorem 4.14].
@@
For the next result, we again rely on the fact that [90, Proposition 4.24] implies that the same recovery can be chosen for all $\mathcal{J}_{n,\Delta_{n,\varepsilon_{n}^{(k)}}}^{(p_{k})}$ with $1\leq k\leq q$. In particular, we need to assume that all $\varepsilon_{n}^{(k)}$ satisfy an appropriate lower bound and that their associated powers $p_{k}$ scale correctly with the dimension of $\Omega$. We then conclude using the subadditivity of $\limsup$.
@@
Summarizing all our previous results and using the subadditivity of $\liminf$ in conjunction with [90, Proposition 4.22], we obtain the following result.
@@
In particular, for the well-posed case, we use [90, Proposition 4.17 and Lemma 4.25] to obtain a uniform bound on the $\mathrm{L}^{2}$-norms of $u_{n}$. Then, uniform and $\mathrm{TL}^{2}$-convergence of a subsequence of $u_{n}$ to some continuous $u$ follows from [90, Proposition 4.13 and Theorem 4.14]. By the uniqueness of the minimizer, Lemma 4.23 and Proposition 2.6, the result follows.
For the ill-posed case, convergence of a subsequence in $\mathrm{TL}^{2}$ to some $u$ follows from [90, Theorem 4.14]. Again, Lemma 4.23 and Proposition 2.6 allow us to conclude. ∎
@@ 5 Numerical Experiments
Since we show that hypergraph learning problem and HOHL behave asymptotically like graph problems in Theorems 3.3, 3.4 and 3.5 (see Figure 2), it is relevant to consider other graph algorithms for fair comparisons: Laplace learning [99], Poisson learning [16], Fractional Laplace (FL) learning [90] with $s=2$ and $s=3$ (only on iris), Weighted Nonlocal Laplacian (WNLL) [75], $p$-Laplace learning [30], Random Walk (RW) [97], Centered Kernel (CK) [59], Sparse LP (SLP) [54] and Properly Weighted Graph-Laplacian [18]. Lastly, we note that no methodical hyperparameter optimization has been performed for the choice of $\varepsilon^{(\ell)}$ and $k^{(\ell)}$. The result highlights are displayed in Tables 3, 4, 5, 6, 7 and 8. The complete results can be found in Appendix 7.

### v2

@@ 4.4 Higher-order hypergraph learning
For our first result, the proof follows from an application of [94, Proposition 4.21]. The key observation is that for any $\varepsilon_{n}$, $p>0$ and $v\in\mathrm{C}^{\infty}(\bar{\Omega})$ with $v(x_{i})=y_{i}$ for $i\leq N$, we can pick $\{(\nu_{n},v_{n})\}_{n=1}^{\infty}=\{(\mu_{n},v|_{\Omega_{n}})\}_{n=1}^{\infty}$ and hence the same recovery sequence for $\mathcal{J}^{(p_{k})}_{n,\Delta_{n,\varepsilon_{n}^{(k)}}}$ with $1\leq k\leq q$. This allows us to use the subadditivity of $\limsup$ to deduce the result.
@@
The next result is shown analogously to Proposition 4.18. In particular, one relies on the compactness result [94, Proposition 4.13]: we only require that our smallest length-scale $\varepsilon_{n}^{(q)}$ satisfies the appropriate upper bound and that the its associated power $p_{q}$ scales correctly with the dimension of $\Omega$. Then, the problem reduces to using the superadditivity of $\liminf$ and [94, Theorem 4.14].
@@
For the next result, we again rely on the fact that [94, Proposition 4.24] implies that the same recovery can be chosen for all $\mathcal{J}_{n,\Delta_{n,\varepsilon_{n}^{(k)}}}^{(p_{k})}$ with $1\leq k\leq q$. In particular, we need to assume that all $\varepsilon_{n}^{(k)}$ satisfy an appropriate lower bound and that their associated powers $p_{k}$ scale correctly with the dimension of $\Omega$. We then conclude using the subadditivity of $\limsup$.
@@
Summarizing all our previous results and using the subadditivity of $\liminf$ in conjunction with [94, Proposition 4.22], we obtain the following result.
@@
In particular, for the well-posed case, we use [94, Proposition 4.17 and Lemma 4.25] to obtain a uniform bound on the $\mathrm{L}^{2}$-norms of $u_{n}$. Then, uniform and $\mathrm{TL}^{2}$-convergence of a subsequence of $u_{n}$ to some continuous $u$ follows from [94, Proposition 4.13 and Theorem 4.14]. By the uniqueness of the minimizer, Lemma 4.26 and Proposition 2.6, the result follows.
For the ill-posed case, convergence of a subsequence in $\mathrm{TL}^{2}$ to some $u$ follows from [94, Theorem 4.14]. Again, Lemma 4.26 and Proposition 2.6 allow us to conclude. ∎
@@ 5 Numerical Experiments
Since we show that hypergraph learning problem and HOHL behave asymptotically like graph problems in Theorems 3.2, 3.3 and 3.4 (see Figure 2), it is relevant to consider other graph algorithms for fair comparisons: Laplace learning [103], Poisson learning [17], Fractional Laplace (FL) learning [94] with $s=2$ and $s=3$ (only on iris), Weighted Nonlocal Laplacian (WNLL) [79], $p$-Laplace learning [32], Random Walk (RW) [101], Centered Kernel (CK) [63], Sparse LP (SLP) [56] and Properly Weighted Graph-Laplacian [19]. Lastly, we note that no methodical hyperparameter optimization has been performed for the choice of $\varepsilon^{(\ell)}$ and $k^{(\ell)}$. The result highlights are displayed in Tables 3, 4, 5, 6, 7 and 8. The complete results can be found in Appendix 7.
@@ 6 Conclusion
Declarations
Data Availability Statement
All data analyzed in this study are publicly available from established open-access repositories, and all data sources are explicitly referenced in the manuscript. No proprietary or restricted data were used.
Funding
AW and AB were supported in part by NSF grant DMS-2152717. MT acknowledges the support of the EPSRC Mathematical and Foundations of Artificial Intelligence Probabilistic AI Hub (grant agreement EP/Y007174/1), the Leverhulme Trust through the Project Award “Robust Learning: Uncertainty Quantification, Sensitivity and Stability” (grant agreement RPG-2024-051) and the NHSBT award 177PATH25 “Harnessing Computational Genomics to Optimise Blood Transfusion Safety and Efficacy”.
Competing interests
The authors have no competing interests to declare that are relevant to the content of this article.